OpenStudy (anonymous):

Choose the equation of the line passing through the point (2, 6) and parallel to y = -3x - 4.

OpenStudy (the_fizicx99):

y= 3x - 4 y-y1 = m(x-x1) plug in your values then solve y-6=-3x - 6 +6 + 6 Y = -3x